When solutions of AgNO₃ and NaOH react, the balanced molecular equation is: 2 AgNO₃(aq) + 2 NaOH(aq) → Ag₂O(s) + 2 NaNO₃(aq) + H₂O(l) How much Ag₂O is produced when 0.200 g of AgNO₃ and 0.200 g of NaOH react?
1
Expert's answer
2020-12-06T13:56:57-0500

Q149255

When solutions of AgNO₃ and NaOH react, the balanced molecular equation is:

2 AgNO₃(aq) + 2 NaOH(aq) → Ag₂O(s) + 2 NaNO₃(aq) + H₂O(l)

How much Ag₂O is produced when 0.200 g of AgNO₃ and 0.200 g of NaOH react?


Solution : The given problem is based on the limiting reactant.


Limiting Reactant : It is the reactant that is completely consumed in the reaction and

which determines the amount of product formed in the chemical reaction.


We are given the mass of both the reactant AgNO3 and NaOH.

We have to find out the Limiting reactant. Once we figure out the Limiting reactant, we can find the mass of product Ag2O produced.


Step 1 : Convert 0.200 g of AgNO₃ and 0.200 g of NaOH to moles .


Molar mass of AgNO₃ = 1 * atomic mass of Ag + 1 * atomic mass of N + 3 * atomic mass of O.

= 1 * 107.87g / mol + 1 * 14.007 g/mol + 3 * 15.999 g/mol

= 107.87g/mol + 14.007 g/mol + 47.997 g/mol

= 169.874 g/mol

moles of AgNO3 = 0.200 g AgNO₃ * 1 mol AgNO₃ / 169.874 g AgNO₃ = 0.001177 mol AgNO₃ .



Molar mass of NaOH = 1 * atomic mass of Na + 1 * atomic mass of O + 1 * atomic mass of H

= 1 * 22.99 g/mol + 1 * 15.999 g/mol + 1 * 1.00794 g/mol

= 22.99 g/mol + 15.999 g/mol + 1.00794 g/mol

= 39.9969 g/mol  


moles of NaOH = 0.200 g NaOH * 1 mol NaOH / 39.9969 g NaOH = 0.005000 mol NaOH .




Step 2 : Calculate the amount of Ag2O formed by each AgNO3 and NaOH, considering that the other reactant is in excess.


2 AgNO₃(aq) + 2 NaOH(aq) → Ag₂O(s) + 2 NaNO₃(aq) + H₂O(l)



0.001177 mol AgNO3 will give = 0.001177 mol AgNO₃ * 1 mol Ag₂O / 2 mol AgNO₃


= 0.0005885 mol Ag₂O


0.005000 mol NaOH will give = 0.005000 mol NaOH . * 1 mol Ag₂O / 2 mol NaOH

= 0.0025 mol Ag₂O


We can see that among AgNO3 and NaOH, AgNO3 is the one which gives less moles of

Ag2O. So AgNO3 is the limiting reactant and the moles of Ag2O formed actually is the one we have calculated from AgNO3 .


Step 3 : Convert 0.0005885 mol Ag2O to grams.


Molar mass of Ag2O = 2 * atomic mass of Ag + 1 * atomic mass of O.

= 2 * 107.87 g/mol + 1 * 15.999 g/mol

= 215.74 g/mol + 15.999 g/mol

= 231.739 g/mol


grams of Ag2O = 0.0005885 mol Ag2O * 231.739 g Ag2O / 1 mol Ag2O


= 0.1364 grams of Ag2O


In the question, we are given the mass of reactants in 3 significant figure, so our

final answer must also be in 3 significant figure.


Hence the mass of Ag2O produced is 0.136 grams.